# An Empirical Study on Cross-lingual Vocabulary Adaptation for Efficient Language Model Inference

### v2

@@ Abstract
Atsuki Yamaguchi1, Aline Villavicencio1,2and Nikolaos Aletras1 1Department of Computer Science, University of Sheffield, United Kingdom 2Department of Computer Science, Institute of Data Science and Artificial Intelligence, University of Exeter, United Kingdom {ayamaguchi1,a.villavicencio,n.aletras}@sheffield.ac.uk
@@ 4.6 Evaluation Metrics
We calculate the average number of prompt tokens per sample for each dataset and tokenizer, and use its relative ratio to each source tokenizer as a proxy for inference speedup following Ahia et al. (2023) and Petrov et al. (2023). We use the average number of prompt tokens rather than the actual inference time because commercial APIs (e.g. OpenAI) often charge users on the basis of the total number of prompt and generated tokens. Note that inference efficiency is independent of the model size.
@@ 7 Recommendations
Our findings suggest that CVA offers substantial inference speedups across tasks, languages and models regardless of the target embedding initialization approach. However, downstream performance is sensitive to the embedding initialization and LoRA fine-tuning. Therefore, we provide the following recommendations to researchers and practitioners:
@@ 8 Conclusion
We use LLMs of various sizes ranging from 1B to 7B, which are far larger than those tested in previous CVA studies. For example, Dobler and de Melo (2023), Liu et al. (2024), and Downey et al. (2023) use XLM-R Conneau et al. (2020) (2.8M). Note that inference efficiency measured by the number of processed (or generated) tokens is not affected by the model size. However, investigating the performance of CVA approaches with larger models would be valuable in future studies.
@@ Acknowledgments
We would like to thank Ahmed Alajrami for proofreading Arabic prompts, and Olga Iakovenko, Constantinos Karouzos, and Huiyin Xue for their comments. We acknowledge IT Services at the University of Sheffield for the provision of services for high-performance computing. AY is supported by the Engineering and Physical Sciences Research Council (EPSRC) [grant number EP/W524360/1] and the Japan Student Services Organization (JASSO) Student Exchange Support Program (Graduate Scholarship for Degree Seeking Students). This work is also partly supported by EPSRC project EP/T02450X/1.
@@ Appendix C Results
C.1 Additional Downstream Results
@@
Although models adapted with Random are competitive to other approaches using in-language prompts (Tables 2 and 10), this is not the case when prompting in English (Table 9). We see a substantial drop in performance, especially in the Japanese, Arabic, and Swahili span tasks with BLOOM-1B. For larger models, Random adversely affects not only span but also performance in sum. For instance, we observe up to 53.0% (German CVA TigerBot-7B) and 60.0% (Japanese CVA TigerBot-7B) performance degradation in zero-shot sum and span, respectively. Our hypothesis is that random initialization can severely impair the ability of the LLM to understand English prompts, which cannot be fully recovered with LAPT due to the exclusion of English data.
@@
C.2 English Downstream Performance
@@
C.3 How helpful is LAPT for LLMs with cross-lingual vocabulary adaptation?

### v3

@@ Abstract
Atsuki Yamaguchi1, Aline Villavicencio1,2,3and Nikolaos Aletras1 1School of Computer Science, University of Sheffield, United Kingdom 2Department of Computer Science, Institute of Data Science and Artificial Intelligence, University of Exeter, United Kingdom 3The Alan Turing Institute, United Kingdom {ayamaguchi1,a.villavicencio,n.aletras}@sheffield.ac.uk
@@ 4.6 Evaluation Metrics
We calculate the average number of prompt tokens per sample for each dataset and tokenizer, and use its relative ratio to each source tokenizer as a proxy for inference speedup following Ahia et al. (2023) and Petrov et al. (2023). We use the average number of prompt tokens rather than the actual inference time because commercial APIs (e.g. OpenAI) often charge users on the basis of the total number of prompt and generated tokens. Note that inference efficiency is independent of the model size. Moreover, previous work Petrov et al. (2023); Hong et al. (2024) has shown a strong correlation between the length of tokenized inputs and actual processing times, i.e. longer input sequence leads to longer processing time.
@@ 7 Recommendations
Our findings suggest that CVA offers substantial inference speedups across tasks, languages and models regardless of the target embedding initialization approach. However, downstream performance is sensitive to the embedding initialization and LoRA fine-tuning. Therefore, we provide the following recommendations to researchers and practitioners [^14] If we do not care about inference efficiency at all, it might make sense to use LAPT instead of CVA methods.:
@@ 8 Conclusion
We use LLMs of various sizes ranging from 1B to 7B, which are far larger than those tested in previous CVA studies. For example, Dobler and de Melo (2023), Liu et al. (2024), and Downey et al. (2023) use XLM-R Conneau et al. (2020) (0.28B). Note that inference efficiency measured by the number of processed (or generated) tokens is not affected by the model size. However, investigating the performance of CVA approaches with larger models would be valuable in future studies.
@@ Acknowledgments
We would like to thank Ahmed Alajrami for proofreading Arabic prompts, and Olga Iakovenko, Constantinos Karouzos, and Huiyin Xue for their valuable feedback. AY is supported by the Engineering and Physical Sciences Research Council (EPSRC) [grant number EP/W524360/1] and the Japan Student Services Organization (JASSO) Student Exchange Support Program (Graduate Scholarship for Degree Seeking Students). This work is also partly supported by the EPSRC project EP/T02450X/1.
@@ Appendix C Results
C.1 Perplexity
Table 9 shows the perplexities of adapted models measured on the language-specific subset of Wikipedia 50K articles.
Model German Japanese Arabic Swahili
BLOOM-1B
Source 45.6 44.7 14.6 45.4
LAPT 22.6 21.0 20.9 55.4
Random 58.5 55.1 53.7 305.1
CLP 75.2 62.6 46.5 190.2
Heuristics 71.5 51.2 46.0 176.5
FOCUS 70.8 50.1 46.1 168.9
CLP+ 75.5 48.8 45.6 185.5
BLOOM-7B
Source 18.7 21.4 9.5 14.9
LAPT 13.6 13.7 11.0 19.4
Random 45.0 54.1 44.5 179.1
CLP 49.1 165.9 30.9 91.8
Heuristics 36.4 43.5 32.0 85.8
FOCUS 36.8 42.3 32.0 82.4
CLP+ 37.0 41.8 29.9 89.7
TigerBot-7B
Source 7.6 8.2 5.1 30.9
LAPT 8.6 10.0 3.0 9.5
Random 116.2 77.8 130.7 636.3
CLP 87.6 39.9 103.0 688.3
Heuristics 47.3 41.4 124.1 605.4
FOCUS 43.9 39.2 81.8 398.6
CLP+ 45.2 38.8 81.6 559.4
Mistral-7B
Source 4.5 8.0 3.7 18.4
LAPT 5.5 9.7 2.9 7.5
Random 78.0 54.4 77.2 587.7
CLP 33.7 41.4 74.5 358.6
Heuristics 34.8 41.4 81.5 413.3
FOCUS 33.4 41.0 66.1 287.4
CLP+ 35.4 40.0 58.1 297.0
Table 9: Perplexity on the language-specific subset of Wikipedia 50K articles. Bold and underlined indicate the best and second-best perplexities across adapted models with the same base model for each language. Note that perplexities are not comparable between models with grey and others due to their difference in vocabulary.
C.2 Additional Downstream Results
@@
Although models adapted with Random are competitive to other approaches using in-language prompts (Tables 2 and 11), this is not the case when prompting in English (Table 10). We see a substantial drop in performance, especially in the Japanese, Arabic, and Swahili span tasks with BLOOM-1B. For larger models, Random adversely affects not only span but also performance in sum. For instance, we observe up to 53.0% (German CVA TigerBot-7B) and 60.0% (Japanese CVA TigerBot-7B) performance degradation in zero-shot sum and span, respectively. Our hypothesis is that random initialization can severely impair the ability of the LLM to understand English prompts, which cannot be fully recovered with LAPT due to the exclusion of English data.
@@
C.3 English Downstream Performance
@@
C.4 How helpful is LAPT for LLMs with cross-lingual vocabulary adaptation?